OpenStudy (anonymous):

Find three consecutive odd integers such that four times the sum of the first and second is 17 more than 7 times the third

OpenStudy (anonymous):

37,39,41

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

let us take 3 consecute ints is x+1,x+3,x+5 then according to given ques the equation we get is 4(x+1+x+3)=17+7(x+5) by soloving this we get x=36 so the three consecutiive integers are 37,39,41